40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 9 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  16084   Sun Apr 25 21:21:02 2021 ranaUpdateCDSSUS simPlant model
  1. I suggest not naming this the LSC model, since it has no LSC stuff.
  2. Also remove all the diagnostic stuff in the plant model. We need nothing except a generic filter Module, like in the SUS controller.
  3. Also, the TF looks kind of weird to me. I would like to see how you derive that eq.
  4. Connect the models and show us some plots of the behavior in physical units using FOTON to make the filters and diaggui/DTT (at first) to make the plots.
  16083   Fri Apr 23 19:26:58 2021 KojiUpdatePSLHEPA speed lowered

I believe that there is an internal setting for the minimum flow, so the flow is not linear ("0%" is not zero), but we should mark this flow speed once you find this is sufficiently low for the locking too.

  16082   Fri Apr 23 18:00:02 2021 gautamUpdatePSLHEPA speed lowered

I will upload some plots later - but in summary, I set the HEPA speed to ~40%. I used (i)IMC transmission RIN, (ii) Arm cavity transmission RIN and (iii) ALS beat noise as 3 diagnostics, to see how noise in various frequency bands for these signals change as a function of the HEPA speed. The MC2T RIN shows elevated noise between 1-10Hz at even the lowest speed I tried, ~20% of the max on each blower. The elevated noise extended to ~50-70 Hz for HEPA speeds >40% of the maximum, and the arm cavity RIN and ALS signals also start to become noisy for speeds >60% of the maximum. So I think 40% is a fine speed to run at - for squeezing measurement we may have to turn off the HEPA for 10mins but for the usual single arm / PRMI / DRMI locking, this should be just fine. For the elevated ALS noise - I'm not sure if the coupling is happening over the top of the enclosure where the fiber bringing light from EX comes close to the HEPA filters, or if it is happening inside the PSL enclosure itself, near the beat mouth - but anyways, at the 40% speed, I don't see any effect on the ALS noise.

I checked with a particle counter at the SW corner of the PSL table (which is the furthest away we can be on the table from the HEPA blowers) after leaving the blowers on for ~30mins and it registered 0 for both 0.3um and 0.5um sized particles (if the blowers are off, the respective numbers are 43 and 9 but I forgot what the units were, and I believe they have to be multiplied by 10). 

I have not yet marked the speed control units yet in case there is some other HEPA science that needs to be done before deciding what is the correct setting. But I think I can get the PRFPMI lock without much issue with this lower speed, which is what I will try later today evening.

Attachment 1: HEPAdiag.pdf
HEPAdiag.pdf
  16081   Fri Apr 23 15:52:19 2021 Ian MacMillanUpdateCDS40m LSC simPlant model

I have attached the framework that I am using for the full system. Plantframework.pdf has the important aspects that I will be changed. Right now I am trying to keep it mostly as is, but I have disconnected the Optic Force Noise and hope to disconnect the Suspension Position Noise. The Optic Force Noise Block is additive to the signal so eliminating it from the system should make it less realistic but simpler. It can be added back easily by reconnecting it.

The next step is adding my plant response, which is simply the transfer function and measurement from the last post. These should be inserted in the place of the red TM_RESP in the model.

The TM_RESP block takes in a vector of dimension 12 and returns a vector of dimension 6. The confusing part is that the block does not seem to do anything. it simply passes the vector through with no changes. I'm not sure why this is the case and I am looking for documentation to explain it but haven't found anything. As to how a 12 vector turns into a 6 vector I am also lost. I will probably just disconnect everything but the x position.

I tried to just throw in my model (see Simple_Plant.pdf) and see what happened but the model would not let me add built-in blocks to the model. This is weird because all the blocks that I am adding are part of the basic library. My guess is that this mode will only accept blocks from the CDL library. I will either need to change my blocks to be made from blocks in the CDL library or maybe I can pass the signal out of the plant framework model then into my model then back to the plant framework model. I think this is just a Matlab thing that I don't know about yet. (Jon probably knows)

I have also attached an image of the controls model for reference. It looks like a mess but I'm sure there is a method. I won't get lost in going through it just assume it works... for now.

The next question I have been asking is how do I show that the system works. When anchal and I made a python simulation of the system, we tested it by seeing the evolution of the degrees of freedom over time given some initial conditions. We could see the pendulum propagating and some of the coupling between the DOFs. This is a fast and dirty way to check if everything is working and should be easy to add. I simply recorded the POS signal and graph it over time. Once we get to a state-space model we can test it by taking the transfer function, but since our plant is literally already just a transfer function there isn't much of a point yet.

Also, I need to add color to my Simple_Plant.pdf model because it looks really boring :(

Attachment 1: Plant_framework.pdf
Plant_framework.pdf
Attachment 2: Simple_Plant.pdf
Simple_Plant.pdf
Attachment 3: Controls.pdf
Controls.pdf
  16080   Thu Apr 22 17:28:34 2021 YehonathanUpdatePSLLaser amplifier

According to the aLIGO 70W amplifier interlock concept the flow rate of the chiller should be between 5 and 40 l/min. The chillers I found in the TCS lab both have around 15 l/min flow rate so we should be fine in that regard.

Assuming that the power consumption of the diode box is ~800W and that the optical output power of the diode is ~ 300W of optical power, the chillers need to be able to remove the remaining power. At room temperature, they both have enough cooling capacity according to their specs.

As for the idea to put the chiller and diode box in the drill room: There are not a lot of options here. The only viable place is the SW corner (attachment 1). I was told this place is used sometimes for liquid nitrogen dewar. Alternatively, if possible, we can move the fire extinguishers to the SW corner and use the NW corner. In that way, we don't have to clear all the junk from the SW corner, as long as the extinguishers are still accessible.

I made a sketch (attachment 2) showing a possible setup for a diode box + chiller rack. The fiber and network cable can go through the hole in the wall that already exists for the N2. It will have to get bigger though (attachment 3). The rack would also need to host some Acromag unit to convert the communication channel of the chiller/flow meter to Ethernet. The Acromag on 1X7 has no spare channels.

The only power socket in the room, to which the drill is connected, is circuit #36 which is connected to panel L in the lab. The breaker's ampacity seems to be 20A if I'm reading the number on the breaker correctly.

 

Attachment 1: 20210422_124940.jpg
20210422_124940.jpg
Attachment 2: DrillRoomSchematics.pdf
DrillRoomSchematics.pdf
Attachment 3: 20210422_125240_1.png
20210422_125240_1.png
  16079   Thu Apr 22 17:04:17 2021 gautamUpdateSUSSettings restored

Indeed, you can make your own snapshot by specifying the channels to snap in a .req file. But what I meant was, we should confirm that all the channels that we modify are already in the existing snapshot files in the autoburt dir. If it isn't, we should consider adding it. I think the whole burt system needs some cleaning up - a single day of burt snapshots occupies ~400MB (!) of disk space, but I think we're recording a ton of channels which don't exist anymore. One day...

Quote:

Your message suggests that we can set burt to start noticing channel changes at home point and create a .req file that can be used to restore later. We'll try to learn how to do that. Right now, we only know how to burt restore using the existing snapshots from the autoburt directory, but they touch more things than we work on, I think. Or can we just always burt restore it to morning time? If yes, what snapshot files should we use?

  16078   Thu Apr 22 15:36:54 2021 AnchalUpdateSUSSettings restored

The mix up was my fault I think. I restored the channels manually instead of using burt restore. Your message suggests that we can set burt to start noticing channel changes at home point and create a .req file that can be used to restore later. We'll try to learn how to do that. Right now, we only know how to burt restore using the existing snapshots from the autoburt directory, but they touch more things than we work on, I think. Or can we just always burt restore it to morning time? If yes, what snapshot files should we use?

  16077   Thu Apr 22 15:34:54 2021 AnchalUpdatePSLPMC transmission

Koji mentioned that the mode of the laser is different for lower diode currents. So that might be the reason why we got less transmission at the low input power but more afterward.

  16076   Thu Apr 22 15:15:26 2021 gautamUpdatePSLPMC transmission

I was a bit surprised by these numbers suggesting the PMC transmission is only 50-60%. I went to the table today and confirmed that it is more like 85% (1.3 W in, 1.1 W transmitted, both numbers from with the FieldMate power meter), as I claimed in 2019. Even being conservative with the power meter errors, I think we can be confident T_PMC will be >80% (modulo any thermal effects with higher power degrading the MM). There isn't any reliable record of what the specs of the PMC mirrors are, but assuming the IO couplers have T=4000ppm and the end mirror has T=500ppm as per Alan's plot, this is consistent with a loss of something like 300ppm loss per mirror - seems very high given the small beam spots, but maybe these mirrors just aren't as high quality as the test masses?

It's kind of unfortunate that we will lose ~20% of the amplifier output through the first filter, but I don't see an easy way to clean these mirrors. It's also not clear to me if there is anything to be gained by attempting a cleaning - isn't the inside of the cavity supposed to be completely isolated from the outside? Maybe some epoxy vaporization events degraded the loss?

Quote:

The transmitted power was ~50-60 mW. (Had to use power meter suspended by hand only.

  16075   Thu Apr 22 14:49:08 2021 gautamUpdateComputer Scripts / Programsrossa added to RTS authorized keys

This is to facilitate running of scripts like the CDS reboot script, mx_stream restart, etc, from rossa, without being pwd prompted every time, whereas previously it was only possible from pianosa. I added the public key of rossa to FB and the RT FE servers. I suppose I could add it to the Acromag servers too, but I haven't yet.

  16074   Thu Apr 22 14:41:55 2021 ChubUpdatePSLNew HEPA speed control

When adjusting the blower speed, give the blower at least 30 seconds to speed up or slow down to the set speed.  The flywheel effect of the big motor armature and blower mass requires time to follow the control current.  Note the taller Flanders HEPA filters.  These and the new intake filters should keep the PSL air clean for a long time!

Quote:

The new HEPA speed controllers are attached at the middle of the HEPA unit (not at the edge of the unit)... (Attachment 1)
You still need a step./stool to touch the knob and need a ladder for a more precise setting.

We still don't know the optimal speed of the nominal IFO operation. For now, the HEPAs are running at the max speed (Attachment 2).
Once we know the optimal setting, we mark the knobs so that we can see them only with the step.

 

Attachment 1: 40M_PSL_HEPA.jpg
40M_PSL_HEPA.jpg
  16073   Thu Apr 22 14:22:39 2021 gautamUpdateSUSSettings restored

The MC / WFS stability seemed off to me. Trending some channels at random, I saw that the MC3 PIT/YAW gains were restored mixed up (PIT was restored to YAW and vice versa) in the last day sometime - I wasn't sure what other settings are off so I did a global burtrestore from the last time I had the interferometer locked since those were settings that at least allow locking (I am not claiming they are optimal).

How are these settings being restored after the suspension optimization? If the burtrestore is randomly mixing up channels, seems like something we should be worried about and look into. I guess it'd also be helpful to make sure we are recording snapshots of all the channels we are changing - I'm not sure if the .req file gets updated automatically / if it really records every EPICS record. It'd be painful to lose some setting because it isn't recorded.


Unconnected to this work - the lights in the BS/PRM chamber were ON, so I turned them OFF. Also unconnected to this work, the summary pages job that updates the "live" plots every half hour seem to be dead again. There is a separate job whose real purpose is to wait for the data from EOD to be transferred to LDAS before filling in the last couple of hours of timeseries data, but seems to me like that is what is covering the entire day now.

Attachment 1: MCdamping.png
MCdamping.png
  16072   Thu Apr 22 12:17:23 2021 Anchal, PacoUpdateSUSMC1 and MC3 Suspension Optimization Summary
MC1 Coil Balancing DC and AC Gains
  POS (DC coil Gain) PIT (DC coil Gain) YAW (DC coil Gain) Coil Output Gains (AC)
UL 0.6613 1 1 0.5885
UR 0.7557 1 -1 0.1636
LL 1.3354 -1 1 1.8348
LR 1.0992 -1 -1 0.5101

Note: The AC gains were measured by keeping output matrix to ideal values of 1s. When optimizing DC gains, the AC gains were uploaded in coil ouput gains.


MC1 Diagonalized input matrix
  UL UR LR LL SIDE
POS 0.1700 0.1125 0.0725 0.1300 0.4416
PIT 0.1229 0.1671 -0.1021 -0.1463 0.1567
YAW 0.2438 -0.1671 -0.2543 0.1566 -0.0216
SIDE 0.0023 0.0010 0.0002 0.0015 0.0360

MC1 Suspension Damping Gains
  Old gains New Gains
SUSPOS 120 270
SUSPIT 60 180
SUSYAW 60 180


MC3 Coil Balancing DC and AC Gains
  POS (DC coil Gain) PIT (DC coil Gain) YAW (DC coil Gain) Coil Output Gains (AC)
UL 1.1034 1 1 0.8554
UR 1.1034 1 -1 -0.9994
LL 0.8845 -1 1 -0.9809
LR 0.8845 -1 -1 1.1434

Note: The AC gains were measured by keeping output matrix to ideal values of 1s. When optimizing DC gains, the AC gains were uploaded in coil ouput gains.


MC3 Input matrix (Unchanged from previous values)
  UL UR LR LL SIDE
POS 0.28799 0.28374 0.21201 0.21626 -0.40599
PIT 2.65780 0.04096 -3.2910 -0.67420 -0.72122
YAW 0.60461 -2.7138 0.01363 3.33200 0.66647
SIDE 0.16601 0.19725 0.10520 0.07397 1.00000

MC3 Suspension Damping Gains
  Old gains New Gains
SUSPOS 200 500
SUSPIT 12 35
SUSYAW 8 12
  16071   Thu Apr 22 08:50:21 2021 AnchalUpdateSUSMC2 Suspension Optimization summary

Yes, during the AC balancing, POS column was set to all 1. This table shows the final values after all the steps. The first 3 columns are DC balancing results when output matrix was changed. While the last column is for AC balancing. During AC balancing, the output matrix was kept to ideal position as you suggested.

Quote:

the POS column should be all 1 for the AC balancing. Where did those non-1 numbers come from?

 

  16070   Thu Apr 22 01:42:38 2021 KojiSummaryElectronicsHV Supply Comparison

New HV power supply from Company 'M' has been delivered. So I decided to compare the noise levels of some HV supplies in the lab. There are three models from companies 'H', 'K', and 'M'.

The noise level was measured with SR785 via Gautam's HP filter with protection diodes.

'H' is a fully analog HV supply and the indicator is analog meters.
'K' is a model with a LCD digital display and numerical keypad.
'M' is a model with a knob and digital displays.

All the models showed that the noise levels increased with increased output voltage.

Among these three, H showed the lowest noise. (<~1uV/rtHz@10Hz and <50nV/rtHz@100Hz) (Attachment 1)

K is quite noisy all over the measured freq range and the level was <50uV/rtHz. Also the PSD has lots of 5Hz harmonics. (Attachment 2)

M has a modest noise level (<~30uV/rtHz@10Hz and <1uV/rtHz@100Hz)except for the noticeable line noise (ripple). (Attachment 3)

The comparison of the three models at 300V is Attachment 4. The other day Gautam and I checked the power spectrum of the HV coil driver with KEPCO and the output noise level of the coil driver was acceptable. So I expect that we will be able to use the HV supply from Company M. Next step is to check the HV driver noise with the model by M used as the supply.

Attachment 1: HV_Supply_PSD_H.pdf
HV_Supply_PSD_H.pdf
Attachment 2: HV_Supply_PSD_K.pdf
HV_Supply_PSD_K.pdf
Attachment 3: HV_Supply_PSD_M.pdf
HV_Supply_PSD_M.pdf
Attachment 4: HV_Supply_PSD.pdf
HV_Supply_PSD.pdf
  16069   Wed Apr 21 19:43:20 2021 KojiUpdatePSLNew HEPA speed control

The new HEPA speed controllers are attached at the middle of the HEPA unit (not at the edge of the unit)... (Attachment 1)
You still need a step./stool to touch the knob and need a ladder for a more precise setting.

We still don't know the optimal speed of the nominal IFO operation. For now, the HEPAs are running at the max speed (Attachment 2).
Once we know the optimal setting, we mark the knobs so that we can see them only with the step.

Attachment 1: P_20210421_193637.jpg
P_20210421_193637.jpg
Attachment 2: P_20210421_193627.jpg
P_20210421_193627.jpg
  16068   Wed Apr 21 19:28:03 2021 AnchalUpdatePSLPSL/IFO recovery

[Anchal, Koji]

Removed the top sheet

  • Opened first from the door side so that any dust would spill outside.
  • Then rolled the sheet inward to meet in the middle.
  • Repeated this twice for the 2 HEPA filters.

Removed the sheets on the table

  • Lifted sheet up making sure the top side face outside always.
  • Rolled it sideways halfway through.
  • Cut down the sheet vertically.
  • Slided the doors to the other side and rolled the remaining half.
  • On the door side, the sheets above the ALS optics were simply lifted off.

Restarting PSL

  • Turned on the HEPAs at the max speed
  • Switched on laser to jsut above the threshold
    • Before the 1st eom, power was 20mW 
    • After the EOM/AOM, 18mW. So about 90% transmission through all polarizing optics.
    • We saw the resonances of the PMC but could not lock it even with highest gain available (30 dB).
  • Increased the input power to PMC to 100mW
    • Locked the PMC at 30dB gain
    • The transmitted power was ~50-60 mW. (Had to use power meter suspended by hand only.
    • The right before the IMC (after the 2nd EOM) 48mW. So none of the alignment was lost.
  • Opened the PSL shutter.
  • We were able to see IMC reflection signal.
  • We were also able to see IMC catching lock as the servo was left ON earlier.
  • Switched off the servo.
  • Decided to increase the power while watching PMC Trans/Refl and IMC REFL
  • Injection diode current to innolight was increased slowly to 2.10A. Saw a mod hopping region aroun 1.8A.
  • We recovered the PMC Trans >0.7 V.
  • PZT was near the edge, so moved by one FSR.
  • The PMC refelction signal is still shown in red at around 48 mV.

Back to control room

  • IMC was locked almost immediately by manually finding the lock while keeping IMC WFS off to preserve the offsets from yesterday.
  • Then switch on IMC WFS. Working good.
  • Then unlocked the servo and switched on IMC Autolocker. Lock was caught immediately.

Decided to start locking the arms

  • The arm transmissions were flashing but at 0.2~0.3 level.
  • Decided to adjust TT1 and TT2 Pitch and Yaw to align the light going into the arms.
  • This made TRY ~0.6 / TRX ~0.8 at the peak of the flashing
  • Locked the arms. (By switching on C1:LSC-MODE_SELECT which engages all servos).
  • Used ASS to align Yarm then align Xarm. Procedure:
    • Sitemap > ASC > c1ass
    • Open striptool to look at progress. ! Scripts YARM > striptool.
    • Switch on ASS. ! More Scripts > ON
    • Wait for the TRY to reach to around 0.97.
    • Freeze the outputs. ! Scripts > Freeze Outputs.
    • Offload the offsets to preserve the output. ! More Scripts > OFFLOAD OFFSETS.
    • Switch off ASS. ! More Scripts > OFF
    • Repeted this for XARM.
  • At the end, both XARM and YARM were locked with TRX ~ 0.97 and TRY ~ 0.96.
  16067   Wed Apr 21 18:49:29 2021 ranaUpdateSUSMC2 Suspension Optimization summary

the POS column should be all 1 for the AC balancing. Where did those non-1 numbers come from?

  16066   Wed Apr 21 15:50:01 2021 Anchal, PacoUpdateSUSMC2 Suspension Optimization summary
MC2 Coil Balancing DC and AC Gains
  POS PIT YAW COIL_GAIN (AC balancing)

UL

1.038 1 1 1.05
UR 1.009 1 -1 -1.05
LL 0.913 -1 1 -1.030
LR 0.915 -1 -1 0.995

MC2 Diagonalized input matrix
  UL UR LR LL SIDE
POS 0.2464 0.2591 0.2676 0.2548 -0.1312
PIT 1.7342 0.7594 -2.494 -1.5192 -0.0905
YAW 1.2672 -2.0309 -0.9625 2.3356 -0.2926
SIDE 0.1243 -0.1512 -0.1691 0.1064 0.9962

MC2 Suspension Gains
  Old gain New Gain
SUSPOS 150 150
SUSPIT 10 30
SUSYAW 10 30

 

  16065   Wed Apr 21 13:10:12 2021 KojiUpdateGeneralPSL HEPA Maintenance

It's probably too late to say but there are/were two boxes. (just for record)

 

  16064   Wed Apr 21 12:56:00 2021 JordanUpdateVACEmpty N2 Tanks

Installed T2 today, and leaked checked the entire line. No issues found. It could have been a bad valve on the tank itself. Monitored T2 pressure for ~2 hours to see if there was any change. All seems ok.

Quote:

When I came into the lab this morning, I noticed that both N2 tanks were empty. I had swapped one on Friday (4-16-21) before I left the lab. Looking at the logs, the right tank (T2) sprung a leak shortly shortly after install. I leak checked the tank coupling after install but did not see a leak. There could a leak further down the line, possibly at the pressure transducer.

The left tank (T1) emptied normally over the weekend, and I quickly swapped the left tank for a full one, and is curently at ~2700 psi. It was my understanding that if both tanks emptied, V1 would close automatically and a mailer would be sent out to the 40m group. I did not receive an email over the weekend, and I checked the Vac status just now and V1 was still open.

I will keep an eye on the tank pressure throughout the day, and will try to leak check the T2 line this afternoon, but someone should check the vacuum interlocks and verify.

 

 

Attachment 1: Screenshot_2021-04-21_12-53-26.png
Screenshot_2021-04-21_12-53-26.png
  16063   Wed Apr 21 11:38:27 2021 Anchal, PacoUpdateSUSMC2 Damping Gains Optimized

We did a step response test with MC2 Suspensoin Damping Gains and optimized them to get <5 oscillations in ringdown.


Procedure:

  • We uploaded the diagonalized input matrix.
  • We uploaded the coil balancing gains at high frequencies found in 16054.
  • We applied Eg2CtQ1 filter module for DC gain balancing foun inf 16055.
  • We set TRAMP to 0 in C1:SUS-MC2_SUSPOS_TRAMP, C1:SUS-MC2_SUSPIT_TRAMP, and C1:SUS-MC2_SUSYAW_TRAMP.
  • We played with offsets to get a good step height. Finally we used:
    • C1:SUS-MC2_SUSPOS_OFFSET: 3000
    • C1:SUS-MC2_SUSPIT_OFFSET: 100
    • C1:SUS-MC2_SUSYAW_OFFSET: 100
  • We looked at channels C1:SUS-MC2_SUSPOS_INMON, C1:SUS-MC2_SUSPIT_INMON, and C1:SUS-MC2_SUSYAW_INMON on a striptool screen to see the step response of the switching on/off of the offsets.
  • We tried to decrease/increase gain to get <5 oscillations during ringdown due to the step inputs.
  • Restored everything back to old values at the end.

Results:

  • Gain in POS was found to be already good. In PIT and YAW we changed the gains from 10 -> 30.
  • Attachment 1 shows the striptool screen when offset was switched ON/Off in POS, PIT and YAW respectively after appling the optimized gains.
  • Attachment 2 shows the same test with old gains for comparison.

In the afternoon, we'll complete doing the above steps for MC1 and MC3. Their coil balancing has not been done on DC so, it is bit non-ideal right now. We'll look into scripting this process as well.

Attachment 1: MC2_DampGainStepTestWithNewGains.png
MC2_DampGainStepTestWithNewGains.png
Attachment 2: MC2_DampGainStepTestWithOldGains.png
MC2_DampGainStepTestWithOldGains.png
  16062   Wed Apr 21 11:09:57 2021 yehonathanUpdatePSLLaser amplifier

I went to the TCS lab to take a look at the chillers lying around. I spotted two chillers:

1. Thermoflex1400 (attachment 1,2). Spec sheet.

2. Polyscience Recirculator 6000 series (attachment 3,4). Manual.

The Thermoflex has various communication ports. The Recirculator doesn't have any communication ports, but it is connected to a flow meter with what seems to be an electronic readout (attachment 5). Manual.

Both chillers have similar capacity ~ 4 gallons/minute. Thermoflex has 2 times more reservoir capacity than the Recirculator.

None of them seem to be Bechkoff-ready.

I guess we can have interlock code handling mixed signals Beckhoff+Non beckhoffs?

Attachment 1: 20210420_171606.jpg
20210420_171606.jpg
Attachment 2: 20210420_171621.jpg
20210420_171621.jpg
Attachment 3: 20210420_171611.jpg
20210420_171611.jpg
Attachment 4: 20210420_171629.jpg
20210420_171629.jpg
Attachment 5: 20210420_171702.jpg
20210420_171702.jpg
  16061   Wed Apr 21 11:01:37 2021 RanaUpdateCDS40m LSC simPlant model

The controller would be in the c1sus model, and connects to the c1sup plant model. So the controller doesn't go in the plant model.

Both the controller and the plant can be modeled using a single filter module in each separate model as you've drawn, but they go in separate models.

 

  16060   Wed Apr 21 10:59:07 2021 ranaSummaryCamerasnote on new GigE cam @ 1064

Note from Stephen on more sensitive Baslers.

  16059   Wed Apr 21 10:03:01 2021 Ian MacMillanUpdateCDS40m LSC simPlant model

So I am stuck on how to add the control block to my model. I am trying to make it as simple as possible with just a simple transfer function for a damped harmonic oscillator and then the control block (see overview.png). 

The transfer function I am using is:

 H(s)=\frac{Q^{-2} f^{-2}s^2+1}{f^{-4}s^4+(Q^{-2} f^{-2}-2f^{-2})s^2+1}

For future generations: To measure the transfer function (to verify that it is doing what it says it is) I am using the discrete transfer function estimator block. To get a quick transfer function estimator Simulink program run ex_discrete_transfer_function_estimator in the Matlab command line. This works well for filters but it was hit or miss using the discrete transfer function. 

The roadblock I am running into right now is that I can't figure out how to add the controller to the model. Not on an interpretation level but in a very straightforward "can I drag it into my model and will it just work" kind of way.

I am also a little confused as to exactly which block would do the controlling. Because I want to just use the x of the pendulum (its position) I think I want to use the suspension controls which come are connected to in the suspension plant model. But where exactly is it and how can I get the model? I can't seem to find it. 

Attachment 1: Overview.png
Overview.png
  16058   Wed Apr 21 05:48:47 2021 ChubUpdateGeneralPSL HEPA Maintenance

Yikes!  That's ONE filter.  I'll get another from storage.

  16057   Wed Apr 21 01:14:03 2021 KojiUpdatePSLPSL Table (sort of) covered / HEPA "chimney"

I also located the (possible) HEPA filters in the lab. (Attachments 1~3)

Oh! This is NO-NO! We can't place anything in front of the mains breakers. (Attachment 2)
I relocated the objects (Attachment 3)

 

Attachment 1: P_20210421_005056.jpg
P_20210421_005056.jpg
Attachment 2: P_20210421_005114.jpg
P_20210421_005114.jpg
Attachment 3: P_20210421_005302.jpg
P_20210421_005302.jpg
  16056   Wed Apr 21 00:08:15 2021 KojiUpdatePSLPSL Table (sort of) covered / HEPA "chimney"

Shutdown Procedure:
PSL Shutter closed / MC Autolocker disabled / PSL mechanical shutter closed / Laser injection current turned to zero / Laser turn off (red button) / Laser key turned off

The laser stat before the shutdown:
- LD Temp A: Set 22.07 (Untouched)
- LD Temp B: Set 21.03(Untouched)
- Laser Injection Current: Dial 9.53, Actual 2.100 -> Dial was moved to zero upon shutting down
- Laser Crystal Temp: Dial 3.34 (untouched)  Set 30.57 Actual 30.60 (Untouched)

PSL Table covering

- Because of the so many cables going up and down, sealing the PSL table with the metalized sheet was not easy. Therefore, the sheets have been just softly laid above the optics. (Attachment 1)
- The largest sheet which covers the east half of the table was taped to the table at the bottom, so that the air from the chimneys (see below) does not come up to the table

- The large dust could come from the opening of the enclosure during the filter replacement. So it was considered to be easier to seal the openings. (Attachment 2)
- Of course, the HEPAs are going to be tested after the maintenance work. It means that vent paths were needed so that the seals do not explode with the pressure (together with dust).
- Thus, the tubes of the sheets are attached to the seals to form "chimneys" for guiding the airflow beneath the table. (Attachment 2/3/4)
- This configuration was not meant to be sufficiently strong for a continuous run of the fans. Long running of the HEPAs may cause the failure of the seal tapes.
  Therefore the HEPA test should be done with a low flow rate and/or a short period of high flow.

- Once the work has been done, all the sheets should be carefully removed without scattering the fallouts onto the optics.

 

Attachment 1: 20210420235324_IMG_0560.jpeg
20210420235324_IMG_0560.jpeg
Attachment 2: 20210420235304_IMG_0559.jpeg
20210420235304_IMG_0559.jpeg
Attachment 3: 20210420235243_IMG_0558.jpeg
20210420235243_IMG_0558.jpeg
Attachment 4: 20210420235344_IMG_0562.jpeg
20210420235344_IMG_0562.jpeg
  16055   Tue Apr 20 18:19:30 2021 AnchalUpdateSUSMC2 coil balanced at DC

Following up from morning's work, I balanced the coils at DC as well. Attachment 1 is screenshot of striptool in which blue and red traces show ASCYAW and ASCPIT outputs when C1:SUS-MC2_LSC_OFFSET was switched by 500 counts. We see very slight disturbance but no real DC offset shown on PIT and YAW due to position step. This data was taken while nominal F2A filter calculated to balance coils at DC was uploaded


I have uploaded the filters on filter banks 7-10 where FM7 is the nominal filter with Q close to 1 and 8-10 are filters with Q 3, 7 and 10 respectively. The transfer function of these filters can be seen in Attachment 2. Note, that the high frequency gain drops a lot when higher Q filters are used.

These filters are designed such that the total DC gain after the application of coil outputs gains for high frequency balancing (as done in morning 16054) balances the coils at DC.


Since I had access to the complete output matrix that balances the coils to less than 1% cross coupling at high frequencies from 16009, I also did a quick test of DC coil balancing with this kind of high frequency balancing. In this case, I uploaded another set of filters which were made at Q close to 1 and gain such that effective DC gain matrix becomes what I got by balancing in the above case. This set of filter also worked as good as the above filters. This completes the proof that we can also use complete matrix for high frequency coil balancing which can be calculated by a script in 20min and works with DC coil balancing as well. In my opinion, this method is more clear and much faster than toggling values in coil output gains where we have only 4 values to optimize 6 cross-coupling parameters. But don't worry, I'm not wasting time on this and will abandon this effort for now, to be taken up in future.


Next up:

  • Tomorrow, we'll finish DC balancing for MC1 and MC3 with the method I practiced today. This should not take much time and should be completed before the meeting.
  • I'll also, calculate and upload the F2A filters for MC1 and MC3.
  • Next, we'll optimize gains in the suspension damping loops by doing step response test (with TRAMP = 0s). We'll look for decaying response (at MC_F, and WFS sensors) with a few oscillations for each step in POS, PIT, and YAW.

Edit Tue Apr 20 21:25:46 2021 :

Corrected the calculation of filters in case of Q different than \large \sqrt{G_{DC}}. There was a bug in the code which I overlooked. I'll correct the filter bank modules tomorrow.


Edit Wed Apr 21 11:06:42 2021 :

I have uploaded the corrected foton filters. Please see attachment 3 for the transfer functions calculated by foton. They match the filters we intended to upload. Only after uploading and closing the foton filter, I realized that the X=7 filter plot (bottom left in attachment 3) does not have dB units on y-axis. It is plotted in linear y-scale (this plot in foton is for phase by default to I guess I forgot to change the scaling when repurposing it for my plot).

Attachment 1: MC2_DC_Coil_Balanced_St.png
MC2_DC_Coil_Balanced_St.png
Attachment 2: IMC_F2A_Params_MC2.pdf
IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf
Attachment 3: UploadedPOS_F2A_Filters.pdf
UploadedPOS_F2A_Filters.pdf
  16054   Tue Apr 20 10:52:49 2021 Anchal, PacoUpdateSUSAC gain coil output balancing for IMC

[Paco, Anchal]

  • We adopted the following procedure to balance the coil output gains using a high-frequency (> 10 Hz) excitation on "C1:SUS-MCX_ASCPIT_EXC", "C1:SUS-MCX_ASCYAW_EXC", and "C1:SUS-MCX_LSC_EXC", where X is one of {1, 2, 3} for the three IMC optics, and the cavity sensors (MC_F, and MC_TRANS);
    1. We load the new input matrix found on March-23rd.
    2. Using awggui, we launch a single 23.17 Hz sine with 500 - 1000 counts amplitude on the aforementioned channels.
      • We are still unable to launch multiple excitations simultaneously through either API (python-awg or dtt-awggui) devilno
    3. Using our built-in hominid neural networks, we look at the "C1:IOO-MC_F", "C1:IOO-MC_TRANS_PIT_IN", and "C1:IOO-MC_TRANS_YAW_IN" exponentially averaging power spectra, on and about the excitation frequency, and identify the amount of cross-coupling going into angular or longitudinal motion depending on the excited degree of freedom.
    4. We step the "C1:SUS-MCX_URCOIL_GAIN", "C1:SUS-MCX_ULCOIL_GAIN", "C1:SUS-MCX_LRCOIL_GAIN", "C1:SUS-MCX_LLCOIL_GAIN" coil output gains by hand in the presence of an excitation (e.g. "LSC") along a given degree of freedom (e.g. along "PIT") to try and minimize the coupling.
    5. We iterate step (4) until we find an optimum gain set, and move on to another optic.

Results

  • For MC2 the optimal gains changed from: [1.0, -1.0, 1.0, -1.0] → [1.05, -1.05, 0.995, -1.03] **
    • Here we were able to first decouple PIT and YAW from a POS excitation almost entirely (see Attachment #1), but weren't as successful in decoupling YAW and POS from PIT, or PIT and POS from YAW excitations (Attachment #2).
  • For MC1 the optimal gains changed from: [1.0, 1.0, 1.0, 1.0] → [0.282, 0.035, 0.302, 2.46] **
    • Here we mostly succeeded in decoupling POS from YAW and PIT excitations (see Attachments #3 - 4).
  • For MC3 the optimal gains changed from: [1.0, -1.0, 1.0, -1.0] → [0.126, -0.123, 0.298, -0.306] **
    • Here the LSC_EXC didn't show up on MC_F (??), and the PIT/YAW excitations decouple by virtue of seemingly low gains, so maybe the optimum is an artifact of the lower coil gains...
    • Plots are to follow up for this one.

** The notation here is [UL, UR, LR, LL]


Attachment 1: POS2PYuncoupled.pdf
POS2PYuncoupled.pdf POS2PYuncoupled.pdf
Attachment 2: PIT2PYuncoupled.pdf
PIT2PYuncoupled.pdf PIT2PYuncoupled.pdf
Attachment 3: MC1YAWexc.pdf
MC1YAWexc.pdf
Attachment 4: MC1PITexc.pdf
MC1PITexc.pdf
  16052   Mon Apr 19 21:54:55 2021 YehonathanUpdateCDSUpdated c1auxey wiring plan

Except for the feed-throughs that require a DB9-M connector I finished wiring and labeling the Acromag, following Jon's updated wiring plan.

We can start testing the differential inputs until the missing connectors arrive.

 

  16051   Mon Apr 19 19:40:54 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Paco, Yehonathan}

We broke the last chunk of ice and cleaned the fridge. We move the fridge back inside and plugged it into the wall. The glues were moved back from the main fridge.

The batteries that were found soaking wet are now somewhat dry and were left on the cabinet drawers for future recycling.

Quote:

{Anchal, Paco, Yehonathan}

We took the glue fridge outside.

 

  16050   Mon Apr 19 13:15:20 2021 Ian MacMillanUpdateCDS40m LSC simPlant model

The x1SUSsim model on the docker was made in a more recent version of Simulink so I updated Matlab (see this)

I updated Matlab to 2021a so now the docker has 2020b and 2021a installed. This should also install Simulink 10.3 for the sus model to open. I used my account to activate it but I can change it over if I get a dedicated license for this. I am not sure what Jon did for the 2020b that he installed.

it is giving me "License error: -9,57" so I guess it didn't work... I will try to just make the model on the 2020b just so I have something.

-----------------------------------------------------------------------------------------------------------------------------

I was able to fix the problem with the activation (using this).

I can now open the x1sussim.mdl file. It is trying to access a BSFM_MASTER Library that it doesn't seem to have. the other files don't seem to have any warnings though. 

the simple suspension model can be found in home/controls/docker-cymac/userapps/c1simpsus.mdl on the docker system. This is where I will put my model. (right now it is just a copied file)

Also using Simulink on the docker is very slow. I think this is either a limit of the x2goclient software or the hardware that the docker is running on.

  16049   Mon Apr 19 12:18:19 2021 Anchal, PacoUpdateSUSTested proposed filters for POS colum in MC2 output matrix

The filters were somewhat successful, how much we can see in attachment 1. The tip about difference between eigenmode basis and cartesian basis was the main thing that helped us take data properly. We still used OSEM data but rotated the output from POS, PIT, YAW to x, theta, phi (cartesian basis where x is also measured as angle projected by suspension length).


Eigenmode basis and Cartesian basis:

  • It is important to understand the difference between these two and what channels/sensors read what.
  • Eigenmode basis as the name suggests is the natural basis for the suspended pendulum.
    • It signifies the motion along three independent and orthogonal modes of motion: POS (longitudinal pendulum oscillation), PIT, and YAW.
    • The position of optic can be written in eigenmode basis as three numbers:
      • POS: Angle made by the center of mass of optic with verticle line from suspension point.
      • PIT: Angle made by the optic face with the suspension wires (this is important to note).
      • YAW: Angle made by optic surface with the nominal plane of suspension wires. (the yaw angle basically).
  • Cartesian basis is the lab reference frame.
    • Here we define three variables that can also represent an optic positioned and orientation:
      • x: Angle made by the center of mass of optic with verticle line from suspension point. (Same as POS)
      • \large \theta: Angle made by the optic surface with absolute verticle (z-axis) in lab frame.
      • \large \phi: Twist of the optic around the z-axis. Same as YAW angle above.
  • We want to apply the feedback gains and filters in eigenmode basis because they are a set of known independent modes. (RXA: NOOO!!!!!! read me elog entry on this topic)
  • Hence, the output from input matrix of suspensions comes out at POS, PIT and YAW in the eigenmode basis.
  • However, the sensors of optic positional, and orientation such at MC_F, wave front sensors and optical levers measure it in lab frame and thus in cartesian basis.
  • Essentially, the \large \theta measured by these sensors is different from the PIT calculated using the OSEM sensor data and is related by:
    • \large \theta = PIT - POS, where PIT and POS both are in radians as defined above.
  • When we optimized the cross-coupling in output matrix at high frequencies using the MC_F and WFS data, we actually optimized it In cartesian basis.
  • The three feedback filters from POS, PIT and YAW which carry data in the eigenmode basis need to be rotated into the cartesian basis in the output matrix before application to the coils.
  • The so-called F2A and A2L filters are essentially doing this rotation.
  • Above the resonant frequencies, the PIT and \large \theta become identical. Hence we want our filters to go to unity

The two filter sets:

  • The filters are named Eg2Ctv1 and Eg2Ctv2 on the POS column of MC2 output matrix.
  • This is to signify that these filters convert the POS, PIT, and YAW basis data (eigenmode basis data) into the cartesian basis (x, theta, phi) in which the output matrix is already optimized at higher frequencies.
  • v1 filter used an ideal output matrix during the calculation of filter as described in 16042 (script at scripts/SUS/OutMatCalc/coilBalanceDC.py).
  • Attachment 2 shows these filter transfer functions.
  • v2 filter use the output matrix optimized to reduce cross-coupling amount cartesian basis modes (MC_F, WFS_PIT and WFS_YAW) in 16009.
  • Attachment 3 shows these filter transfer funcitons.
  • Because of this, the v2 filter is different among right and left coils as well. We do see in Attachment 1 that this version of filter helps in reducing POS->YAW coupling too.

Test procedure:

  • We uploaded both the diagonalized input matrix and the diagonalized output matrix as calculated earlier.
  • We measured channels C1:SUS-MC2_SUSPOS_IN1_DQ, C1:SUS-MC2_SUSPIT_IN1_DQ, and C1:SUS-MC2_SUSYAW_IN1_DQ throughout this test.
  • These channels give output in an eigenmode basis (POS, PIT, and YAW) and the rows of the input matrix have some arbitrary normalization.
  • We normalize these channels to have same input matrix normalization as would be for ideal matrix (2 in each row).
  • Then, assuming the UL_SENS, UR_SENS, LR_SENS, and LL_SENS channels that come at input of the input matrix are calibrated in units of um, we calculate the cartesian angles x, theta, phi. for this calculation, we used the distance between coils as 49.4 mm (got it from Koji) and length of suspension as 0.2489 m and offset of suspension points from COM, b = 0.9 mm.
  • Now that we have true measures of angles in cartesian basis, we can use them to understand the effect on cross coupling from the filters we used.
  • PSL shutter is closed and autolocker is disabled. During all data measurements, we switched of suspension damping loops. This would ensure that our low frequency excitation survives for measurement at the measurement channels.
  • We first took reference data with no excitation and no filters for getting a baseline on each channel (dotted curves in Attachment 1).
  • We then send excitation of 0.03 Hz with 500 counts amplitude at C1:SUS-MC2_LSC_EXC and switched on LSC output.
  • One set of data is taken with no filters active (dashed curve in attachment 1).
  • Then two sets of data are taken with the two filters. Each data set was of 500s in length.
  • Welch function is used to take the PSD of data with bin widht of  0.01Hz and 9 averages.

Results:

  • Filter v1 was the most successful in reducing \large x \rightarrow \theta coupling by factor of 17.5.
    • The reduction in \large x \rightarrow \phi coupling was less. By a factor of 1.4.
  • Filter v2 was worse but still did a reduction of \large x \rightarrow \theta coupling by factor of 7.8.
    • The reduction in \large x \rightarrow \phi coupling was better. By a factor of 3.3.

Next, filters in PIT columns too

  • We do have filters calculated for PIT as well.
  • Now that we know how to test these properly, we can test them tomorrow fairly quickly.
  • For the YAW column though, the filters would probably just undo the output matrix optimization as they are derived from ideal transfer function models and ideally there is no coupling between YAW and other DOFs. So maybe, we should skip putting these on.
Attachment 1: CrossCoupleTestForEgToCtFilters.pdf
CrossCoupleTestForEgToCtFilters.pdf
Attachment 2: uncFilters.pdf
uncFilters.pdf uncFilters.pdf uncFilters.pdf
Attachment 3: uncFilters_v2.pdf
uncFilters_v2.pdf uncFilters_v2.pdf uncFilters_v2.pdf
  16048   Mon Apr 19 10:52:27 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Anchal, Paco, Yehonathan}

We took the glue fridge outside.

  16047   Mon Apr 19 09:17:51 2021 JordanUpdateVACEmpty N2 Tanks

When I came into the lab this morning, I noticed that both N2 tanks were empty. I had swapped one on Friday (4-16-21) before I left the lab. Looking at the logs, the right tank (T2) sprung a leak shortly shortly after install. I leak checked the tank coupling after install but did not see a leak. There could a leak further down the line, possibly at the pressure transducer.

The left tank (T1) emptied normally over the weekend, and I quickly swapped the left tank for a full one, and is curently at ~2700 psi. It was my understanding that if both tanks emptied, V1 would close automatically and a mailer would be sent out to the 40m group. I did not receive an email over the weekend, and I checked the Vac status just now and V1 was still open.

I will keep an eye on the tank pressure throughout the day, and will try to leak check the T2 line this afternoon, but someone should check the vacuum interlocks and verify.

 

Attachment 1: N2_Pressure.PNG
N2_Pressure.PNG
  16046   Sun Apr 18 21:29:55 2021 ranaUpdatePSLLaser amplifier
  • Ideally, we put the chiller outside of the interferometer area. The PSL chiller used to be in the control room near the door by IMC REFL. We could also put it in the drill press room.
  • Once we figure out a couple of places where the Diode Box can go, we can ask facilities to make the appropriate power connections. They will have to eval the situation to figure out if the main power to the lab needs to be shut down.
  • Can we put the laser diode box in the drill press room too? Then the hoses can be short. Perhaps less EMI getting into our sensitive places.
  16045   Fri Apr 16 19:07:31 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

There is still a huge chunk of unmelted ice in the fridge. I moved the content of that fridge in the main fridge and put "do not eat" warning signs.

I returned the fridge to the lab and plugged it back in to prevent flooding.

Defrosting will have to continue on Monday.

  16044   Fri Apr 16 18:21:36 2021 YehonathanUpdatePSLLaser amplifier

I surveyed a bit the 1X1/2 area to plan for the installation of the laser amplifier.

There is a vacancy at the bottom of 1X2 (attachment 1). I measured the dimensions of the diode box (DB) and it should fit. The optical fiber bundle is 75m long and should reach the amplifier head on the table easily.

According to the specs, the maximum power consumption of the DB is 800W (typically 600W), it should probably have its own circuit breaker. It can easily draw more than a few amps. The rack power strips are connected to this 4 socket box (attachment 2), is this just another power strip? It is connected to a circuit breaker with a 30A rating. How do we proceed from here?

In any case, we will need at least 2 meters of power cable.

I also tried to find a suitable place for a water chiller. A few suggestions are in the attachments. Basically either between the electronics shelves and the small rack next to 1X2 or next to the small rack close to the optical table. Maybe put it where the ladder sits and find another place for the ladder. Other options?

We would also need a windows machine running the Beckhoff software. The idea is that all the different laser components (DB, chillers, interlocks, switches) are connected to the EtherCat (over the ethernet infrastructure) so that the Beckhoff code can recognize a failure and switch off everything.

The things that are monitored:

1. Is the NPRO on?

2. Is the flow rate from the chillers enough?

3. Is the temperature of the diodes in the normal range?

4. Is one of the interlocks open?

5. Was one of the emergency buttons pushed?

6. Was the key switch on the DB turned to OFF?

The DB is EtherCat ready but the rest of the signals need to be interfaced somehow. Do we have to buy these EtherCAT terminals?

 

 

Attachment 1: 20210416_143642.jpg
20210416_143642.jpg
Attachment 2: 20210416_145408.jpg
20210416_145408.jpg
Attachment 3: 20210416_145448.jpg
20210416_145448.jpg
Attachment 4: 20210416_181324.jpg
20210416_181324.jpg
  16043   Fri Apr 16 15:47:58 2021 ranaUpdateSUSTested proposed filters for POS colum in MC2 output matrix

Looks mostly right, but you used the OSEM sensors as readbacks. We are diagonalizing using the cavity sensors. Using the diagonalized input matrix is also good since that will reduce the cross-coupling due to the damping loops.

Its sort of a subtle issue:

  1. The sensors are are diagonalized into the eigenmode basis, not the Cartesian basis of the mirror motion.
  2. What the cavity cares about is the Cartesian motion.
  3. Q1: in the model, how much Cartesian pitch motion is there at the POS eigenfrequency in the free-swinging case?
  4. Q2: should we somehow diag the input matrix into the Cartesian basis?
  5. Q3: if so, How?
  6. Q4: and why?
  16042   Fri Apr 16 11:36:36 2021 Anchal, PacoUpdateSUSTested proposed filters for POS colum in MC2 output matrix

We tried two sets of filters on the output matrix POS column in MC2. Both versions failed. Following are some details.


How test was done:

  • PSL shutter was closed and autolocker was switch off.
  • Turned off damping on POS, PIT, and YAW using C1:SUS-MC2_SUSPOS_SW2, C1:SUS-MC2_SUSPIT_SW2, and C1:SUS-MC2_SUSYAW_SW2.
  • Reference data was taken with no excitation to get relative increase at excitation.
  • Channels C1:SUS-MC2_SUSPIT_IN1, C1:SUS-MC2_SUSPOS_IN1, and C1:SUS-MC2_SUSYAW_IN1.
  • Frist we sent an excitation through LOCKIN1 at 0.11 Hz and 500 counts amplitude.
  • LOCKIN column in MC2 output matrix was kept identical to POS column, so all ones.
  • This formed our reference data set when no filters were used. Attachment 1.
  • Note that the peak at 0.03 Hz is due to LOCKIN2 that was left switched on due to autolocker.
  • Then the calculated filters were loaded using foton. Procedure:
    • Right click on filter bank med. Got to Execute-> Foton.
    • Go to File and uncheck 'Read Only'.
    • Find the filter module name in Module drop down.
    • Select an empty module section in Sections.
    • Write a name for the filter. We used DCcoupF2A and DCcouF2A2 for the two version respectively.
    • Paste the zpk foton format in Command.
    • Check with Bode plot if these are correct filters. Then click on Save. It will take about 30s to become responsive again.
    • GO back to filter bank medm screen and click on 'Load Coefficients'. This should start displaying your new filter module.
    • To switch on the module, click on the button below its name.
  • Once fitlers were loaded, we realized we can not use the LOCKIn to excite anymore as it comes as separate excitation.
  • So we used awggui to excite C1:SUS-MC2_LSCEXC at 0.11 Hz and 500 counts.
  • Then we retook the data and checked if the peaks are visible on PIT and YAW channels and how high they are.

Filer version 1

  • This was calculated by starting from ideal output matrix elements as they are currently loaded. All 1's for POS and so on.
  • The calculations were done in scripts/SUS/OutMatCalc/coilBalanceDC.py.
  • This file uses a state space model of the suspension and calculated the cross-coupling. Then the cross coupling is inverted and applied to the current output matrix elements to get correction DC gains.
  • These corrected DC gains are then used to create the filters as described in last post.
  • Attachment 2 shows the filter transfer functions and Attachment 3 shows the test results. Failed :(.
  • There was practivally no change in cross coupling that we can see.

Filter version 2:

  • In this version we used the output matrix optimized at high frequencies earlier (16009).
  • While testing this version, we also uploaded this optimised output amtrix at high frequency.
  • In this test, we realized the LOCKIN2 was on and switched it off manually. All excitations were done through awggui.
  • Attachment 4 shows the filter transfer functions and Attachment 5 shows the test results. Failed :(.
  • There was again practivally no change in cross coupling that we can see.

Forgot to upload new MC2 input matrix:

  • In hindsight, we should have uploaded our diagonalized suspension input matrix in MC2.
  • Without it, there was cross-coupling the in the sensor data to begin with.
  • But this can only be part of the reason why all our filters failed miserably.
  • Because the output matrix was not diagonalized earlier but it was not so bad. Onyl a fresh test can tell if it was the culprit.
Attachment 1: 20210416_MC2DCcoilBalancingNoFilters.pdf
20210416_MC2DCcoilBalancingNoFilters.pdf
Attachment 2: uncFilters.pdf
uncFilters.pdf uncFilters.pdf uncFilters.pdf
Attachment 3: 20210416_MC2DCcoilBalancingWithFilters.pdf
20210416_MC2DCcoilBalancingWithFilters.pdf
Attachment 4: uncFilters_v2.pdf
uncFilters_v2.pdf uncFilters_v2.pdf uncFilters_v2.pdf
Attachment 5: 20210416_MC2DCcoilBalancingWithFilters_v2.pdf
20210416_MC2DCcoilBalancingWithFilters_v2.pdf
  16041   Fri Apr 16 11:31:00 2021 ranaUpdateelogelog stuck ~10 AM today

found it unresponsive. Restarted fine using procedure documented in wiki

  16040   Fri Apr 16 10:58:16 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Paco, Anchal, Yehonathan}

We emptied the fridge and moved the amplifier equipment on top of the amplifier crate. We unplugged the freezer and moved it out of the lab to defrost (attachment).

Quote:

I was looking at the laser head/amp and somehow decided to open the glue freezer. And it was stuck. I've managed to open it but the upper room was completely frozen.
Some of the batteries were embedded in a block of ice. I think we should throw them out.

Can the person who comes in the morning work on defrosting?

- Coordinate with Yehonathan and move the amps and the wooden crate so that you can move the freezer.

- Remove the contents to somewhere (it's OK to be room temp for a while)

- Unplug the freezer

- Leave the freezer outside with the door open. After a while, the ice will fall without care.

- At the end of the day, move it back to the lab. Continue defrosting the other day if the ice remains.

 

Attachment 1: 20210416_105048.jpg
20210416_105048.jpg
  16039   Fri Apr 16 00:21:52 2021 KojiUpdateGeneralGlue Freezer completely frozen

I was looking at the laser head/amp and somehow decided to open the glue freezer. And it was stuck. I've managed to open it but the upper room was completely frozen.
Some of the batteries were embedded in a block of ice. I think we should throw them out.

Can the person who comes in the morning work on defrosting?

- Coordinate with Yehonathan and move the amps and the wooden crate so that you can move the freezer.

- Remove the contents to somewhere (it's OK to be room temp for a while)

- Unplug the freezer

- Leave the freezer outside with the door open. After a while, the ice will fall without care.

- At the end of the day, move it back to the lab. Continue defrosting the other day if the ice remains.

Attachment 1: P_20210416_000906.jpg
P_20210416_000906.jpg
Attachment 2: P_20210416_000850.jpg
P_20210416_000850.jpg
  16038   Thu Apr 15 18:33:39 2021 KojiUpdateCDSMore 16bit ADC adapter boards

We received 10x 16bit ADC adapter boards from Todd. S2100687~S2100696

The number of soldered resistors seems to be less than that on the schematics. They are related to duotone, so check if it's OK upon use.

Attachment 1: P_20210415_183139_1.jpg
P_20210415_183139_1.jpg
  16037   Thu Apr 15 17:24:08 2021 JonUpdateCDSUpdated c1auxey wiring plan

I've updated the c1auxey wiring plan for compatibility with the new suspension electronics. Specifically it is based on wiring schematics for the new HAM-A coil driver (D1100117), satellite amplifier (D1002818), and HV bias driver (D1900163).

Changes:

  • The PDMon, VMon, CoilEnable, and BiasAdj channels all move from DB37 to various DB9 breakout boards.
  • The DB9 cables (x2) connecting the CoilEnable channels to the coil drivers must be spliced with the dewhitening switching signals from the RTS.
  • As suggested, I added five new BI channels to monitor the state of the CoilEnable switches. For lack of a better name, they follow the naming convention C1:SUS-ETMY_xx_ENABLEMon.

@Yehonathan can proceed with wiring the chassis.

Quote:

I finished prewiring the new c1auxey Acromag chassis (see attached pictures). I connected all grounds to the DIN rail to save some wiring. The power switches and LEDs work as expected.

I configured the DAQ modules using the old windows machine. I configured the gateway to be 192.168.114.1. The host machine still needs to be setup.

Next, the feedthroughs need to be wired and the channels need to be bench tested.

Attachment 1: C1AUXEY_Chassis_Feedthroughs_-_By_Connector.pdf
C1AUXEY_Chassis_Feedthroughs_-_By_Connector.pdf
  16036   Thu Apr 15 15:54:46 2021 gautamUpdateIOOWaveplate commissioning - hardware installed

[jordan, gautam]

We did the following this afternoon.

  1. Disconnected the cable from the unused (and possibly not working) RefCav heater power supply, and removed said PS from 1X1. There was insufficient space to install the ESP300 controller elsewhere. I have stored the power supply along the east arm under the beamtube, approximately directly opposite the RFPD cabinet.
  2. Installed the ESP 300 - conveniently, the HP DCPS was already sitting on some rails and so we didn't need to add any.
  3. Ran a long D25-D25 cable from the ESP300 to the NE corner area of the PSL enclosure. The ends of the cable are labelled as "ESP end" and "Waveplate end". The HEPA was turned on for the duration we had the enclosure open, and I have now turned it off.
  4. Connected the waveplate to this cable. Also re-connected the ESP300 to the c1psl supermicro host via the USB-RS232 adapter cable.

The IMC stayed locked throughout our work, and judging by the CDS overview screen, we don't seem to have done any lasting damage, but I will run more tests. Note that the waveplate isn't yet installed in the beam path - I may do this later today evening depending on lab activity, but for now, it is just sitting on the lower shelf inside the PSL enclosure. I will post some photos later.

Quote:
 

So this system is ready to be installed once Jordan and I find some time to lay out cabling + install the ESP300 controller in a rack.


Update: The waveplate was installed. I gave it a couple of rounds of cleaning by first contact, and visually, it looked good to me. More photos uploaded. I also made some minor improvements to the MEDM screen, and setup the communication script with the ESP300 to run as a systemd service on c1psl. Let's see how stable things are... I think the philosophy at the sites is to calibrate the waveplate rotation angle in terms of power units, but i'm not sure how the unit we have performs in terms of backlash error. We can do a trial by requesting ~100 "random" angles, monitoring the power in s- and p-polatizations, and then quanitfying the error between requested and realized angles, but I haven't done this yet. I also haven't added these channels to the set recorded to frames / to the burt snapshot - do we want to record these channels long term?

  16035   Thu Apr 15 11:41:43 2021 AnchalUpdateSUSProposed filters for output matrix aka F2A aka F2P

Here' s aquick update before we leave for lunch. We have managed to calculate some filter that would go on the POS column in MC2 output matrix filter banks aka F2A aka F2P filters. In the afternoon if we can come and work on the IMC, we'll try to load them on the output matrix. We have never done that so it might take some time for us to understand on how to do that. Attached is the bode plot for these proposed filters. Let us know if you have any comments.

Attachment 1: MC2propPOSfb.pdf
MC2propPOSfb.pdf
  16034   Thu Apr 15 09:46:24 2021 YehonathanUpdatePSLLaser amplifier

Some more relevant documents provided by Matt:

Phase III:70W amplifier integration at LIGO

70W amplifier External Shutter

aLIGO PSL high power attenuator

 

ELOG V3.1.3-